OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -3 and contains the point (1, 1).

OpenStudy (anonymous):

Form of the problem will be y=mx+b. We know m: \[y=-3x+b\] Plug in the x and y values: \[1=-3(1)+b\] \[1=-3+b\] \[b=4\] Place the known value of b back into orginal formula: \[y=-3x+4\]